You are doing an experiment, trying to determine which has more effect on potential energy, height or mass. You take two cupcakes. One cupcake you add mass to, and drop from 3 meters. The other you don’t add mass to and drop from 1 meter. Which is more important in potential energy, height or mass? Show why in the equation.
1
Expert's answer
2020-03-26T12:18:24-0400

We know that the potential energy of the body which have some mass m and gravitational acceleration is g. If we are taking that object to certain height, then there will be energy stored in the object that will equal =mgh

where h is the height.

So, potential energy of the object is depending on the mass of the object and height.

gravitational acceleration of the object is constant that is 9.8 m/sec^2

As per the given question,

There is two cup, in which we can add mass

Both the cup have the same mass, let that is m

But in the first cut we can add some, let the net mass of the cup=m+M

Where, M is the external added mass.

It is at height h=3m from the ground.

So, Potential energy

"=(M+m)gh"

"=(M+m)9.8\times 3=29.4(m+M)J"

Now for the second case,

mass of the cup =m and released from the height h=1m

so potential energy =mgh

=9.8mJ